For procedure 4, suppose you measure the diameter of a wire to be 0.499mm and the voltage across the ends to be 35.7mV. What is the resistivity of the wire?

Enter your response to three significant figures.

Given:

Length (L) = 24 cm Current (I) = 1A

I am most concerned with sig figs and decimals. I have it solved for 2.91e-9 but only have one more chance to input correctly.
Answer #1

Using Ohm's law:

V = i*R

R = resistance, which is given by:

R = rho*L/A

rho = resistivity of wire = ?

Area of wire = pi*d^2/4

d = diameter = 0.499 mm = 0.499*10^-3 m

L = length of wire = 24 cm = 0.24 m

So,

V = i*rho*L/(pi*d^2/4)

rho = V*pi*d^2/(4*i*L)

Using given values:

rho = 35.7*10^-3*pi*(0.499*10^-3)^2/(4*1*0.24)

rho = 2.91e-8 ohm-m

Please check your calculations again final answer should be 2.91e-8, let me know if it works

Please Upvote.